1/5 x + 3/2 = 3/10
Answer by waynest(281)

10(1/5)x + 10(3/2) = 10(3/10)
2x + 15 = 3
2x + 15 - 15 = 3 -15
2x = -12
2x/2 = -12/2
x = -6
check
1/5(-6) + 3/2 = 3/10
-6/5 + 3/2 = 3/10
make all over 10
-12/10 + 15/10 = 3/10
3/10 = 3/10